# Ano : Faster is Better in Noisy Landscapes

### v1

@@ 6.2 Transformers
All runs finetune the public bert-base-uncased checkpoint (Devlin et al., 2019) with max sequence length 128, batch size 32, weight decay 0.01, linear schedule with 10% warmup, mixed precision, and 3 epochs (5 for small datasets: CoLA, MRPC, RTE, STS-B). Each configuration is repeated over 5 seeds.
@@
Baseline
AdamW $\mathbf{59.40}\pm\mathbf{1.67}$ $\mathbf{84.62}\pm\mathbf{0.10}$ $88.06\pm 0.82$ $\mathbf{91.60}\pm\mathbf{0.15}$ $89.64\pm 0.10$ $66.67\pm 1.59$ $92.73\pm 0.46$ $88.44\pm 0.27$ $82.64$
@@
Table 3: Average performance (mean $\pm$ 95% CI) of different optimizers on GLUE benchmark tasks (best per column in bold).
@@ 6.3 Deep Reinforcement learning
Reinforcement learning (RL) poses unique challenges, including high gradient variance and the non-stationarity of the environment, both of which can impact the performance of optimization algorithms. Consequently, optimizers that demonstrate strong performance in domains such as natural language processing may behave quite differently in RL settings (Henderson et al., 2019; François-Lavet et al., 2018). This is precisely the domain where Ano is designed to achieve its most significant gains. To evaluate our algorithms, we adopt the MuJoCo suite from the Gymnasium framework developed by the Farama Foundation (Todorov et al., 2012; Towers et al., 2024).
@@
In this section, we employ the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018). We reuse the standard SAC hyperparameter (full list in Appendix E.1 - Tab 11), as reported in the original work and subsequent studies and only vary the optimizer for actor, critics, and temperature. For each optimizer, we run 10 seeds, with 1M steps. We report below the average mean score over the 20 final runs at the end of the training.
@@
Baseline
Adam $9477.97\pm 1003.20$ $4161.08\pm 515.06$ $\mathbf{5247.63}\pm\mathbf{105.87}$ $4276.75\pm 238.68$ $2922.12\pm 284.42$ $2.6$ $90.01$
Adan $8796.97\pm 955.18$ $3294.58\pm 433.50$ $5116.14\pm 125.19$ $3698.10\pm 572.68$ $2509.11\pm 324.34$ $4.8$ $79.87$
Lion $7406.13\pm 2772.84$ $3491.82\pm 579.52$ $85.58\pm 28.33$ $4768.21\pm 201.84$ $703.34\pm 657.37$ $4.6$ $50.72$
Grams $6457.31\pm 879.37$ $3031.85\pm 478.99$ $5181.55\pm 105.09$ $3386.75\pm 896.29$ $2606.07\pm 579.75$ $4.8$ $74.25$
Ano (Ours) $\mathbf{10851.8}\pm\mathbf{518.67}$ $\mathbf{5247.02}\pm\mathbf{653.98}$ $5190.86\pm 136.30$ $\mathbf{5119.08}\pm\mathbf{420.88}$ $\mathbf{2926.39}\pm\mathbf{308.72}$ $\mathbf{1.2}$ $\mathbf{99.78}$
Anolog (Ours) $10389.5\pm 1041.29$ $4598.67\pm 727.16$ $5161.95\pm 268.91$ $4748.18\pm 593.57$ $2585.67\pm 238.48$ $3.0$ $92.57$
Tuned
Adam $9874.24\pm 735.62$ $4110.11\pm 692.96$ $\mathbf{5308.70}\pm\mathbf{195.70}$ $4369.92\pm 326.27$ $\mathbf{3143.48}\pm\mathbf{233.52}$ $2.6$ $90.43$
Adan $10133.64\pm 849.35$ $5222.12\pm 218.59$ $5034.40\pm 95.83$ $\mathbf{4981.25}\pm\mathbf{184.94}$ $2971.66\pm 367.72$ $2.4$ $95.35$
Lion $7351.97\pm 3058.09$ $2826.41\pm 1698.65$ $86.81\pm 24.43$ $4496.17\pm 519.54$ $430.78\pm 260.73$ $4.6$ $44.74$
Grams $8522.14\pm 1648.23$ $4187.69\pm 707.71$ $5116.82\pm 135.39$ $4204.40\pm 836.06$ $2870.06\pm 740.79$ $3.6$ $85.18$
Ano (Ours) $\mathbf{11197.73}\pm\mathbf{278.11}$ $\mathbf{5389.34}\pm\mathbf{353.92}$ $5071.59\pm 138.17$ $4810.42\pm 356.80$ $2972.84\pm 418.39$ $\mathbf{1.8}$ $\mathbf{97.34}$
Table 4: Comparison of the average performance (± 95% CI) of different optimizers across environments.
@@
As summarized in Table 4, Ano performs favorably compared to Adam and other baselines across the MuJoCo tasks. On average, it achieves a +10% improvement in normalized score [^2] The normalized average is obtained by linearly rescaling each score between the minimum and maximum values observed across optimizers, followed by averaging. Complete normalized results are reported in Table 8., both under default and tuned hyperparameters. Without tuning, Ano ranks first in 4 out of 5 tasks; with tuning, it remains the top optimizer in 2 out of 5 tasks. Although not always the best performer, Ano consistently ranks among the strongest optimizers, with its scores typically within or close to the 95% confidence intervals of the best baselines. Figure 2 shows that Ano converges faster, reaching Adam’s final performance in roughly 70% of the training steps. In contrast, Lion and Grams exhibit pronounced instability on high-variance tasks such as Humanoid, while Anolog, our ablated variant—remains competitive but consistently lags behind Ano. To further address potential concerns about hyperparameter tuning, we evaluated sensitivity to the learning rate and momentum coefficients on HalfCheetah (Figure 3) on a 100k steps HalfCheetah proxy. Ano exhibits markedly lower sensitivity than Adam on learning rate and betas, suggesting that its advantage is not merely due to more favorable hyperparameters.
@@
To assess the generality of our findings, we additionally evaluate a discrete-action variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017) on the Atari Learning Environment (ALE) (Bellemare et al., 2013). For compute efficiency, we report results on the Atari-5 subset proposed by Aitchison et al. (Aitchison et al., 2022), which explains $98.4\%$ of the variance in full-suite performance ($R^{2}{=}0.984$). We use the reference PPO implementation from CleanRL (Huang et al., 2022) and keep its default network and optimization hyperparameters (full list in Appendix E.1 - tab 12). Environments are instantiated via EnvPool (Weng et al., 2022) for fast batched simulation. Observations are resized to $84{\times}84$, converted to grayscale, and stacks of 4 consecutive frames are fed to the agent. We apply action repeat of 4, up to 30 no-op actions at reset, the ALE sticky-action protocol with repeat probability $0.25$ (Machado et al., 2017), and FireReset where required. We used the full action set. During training, rewards are clipped to [1,1][-1,1]; evaluation uses unclipped rewards. We train for 1010M agent steps (with action repeat 4, i.e., $\approx 40$M ALE frames), evaluating every 200200k steps on 50 runs. Each checkpoint is evaluated with the same wrappers as training (except reward clipping). We report the average final score on the last evaluation. The normalization average is computed in the same ways that previous part [^3] We normalized each score by environment by the theoretically minimum (-18 for DoubleDunk, 0 for the others) and the maximum score obtained by the baselines, we then compute the mean average normalized score..
@@
Baseline
Adam $7656.00\pm 812.20$ $\mathbf{-0.91}\pm 0.34$ $733.92\pm 107.34$ $\mathbf{3516.64}\pm 475.14$ $4322.00\pm 254.93$ $2.2$ $92.40$
Adan $6408.00\pm 1525.14$ $0.98\pm 0.21$ $633.80\pm 44.93$ $2033.76\pm 85.42$ $2100.60\pm 951.20$ $4.8$ $70.22$
Lion $1420.00\pm 256.88$ $1.96\pm 1.50$ $564.44\pm 195.39$ $2477.12\pm 490.65$ $3680.80\pm 1057.34$ $5.2$ $63.83$
Grams $\mathbf{8396.00}\pm\mathbf{1565.13}$ $1.07\pm 0.55$ $644.64\pm 48.73$ $2133.76\pm 234.64$ $3646.70\pm 1139.41$ $3.8$ $81.41$
Ano (Ours) $8280.00\pm 916.31$ $1.03\pm 0.21$ $\mathbf{806.56}\pm\mathbf{114.49}$ $2808.68\pm 465.32$ $5122.12\pm 783.72$ $\mathbf{2.2}$ $\mathbf{94.48}$
Anolog (Ours) $7620.00\pm 1225.14$ $1.07\pm 0.26$ $743.00\pm 191.30$ $2815.50\pm 430.55$ $\mathbf{5413.75}\pm\mathbf{1873.65}$ $2.8$ $92.40$
Tuned
Adam $7880.00\pm 1218.48$ $\mathbf{-0.93}\pm\mathbf{0.15}$ $663.87\pm 178.56$ $2964.16\pm 601.40$ $4222.90\pm 256.83$ $2.6$ $90.51$
Adan $4672.00\pm 2542.63$ $1.08\pm 0.27$ $693.96\pm 87.13$ $\mathbf{3020.48}\pm\mathbf{419.22}$ $4540.20\pm 996.34$ $2.6$ $86.03$
Lion $1296.00\pm 164.16$ $2.73\pm 1.92$ $550.96\pm 134.29$ $2554.80\pm 546.74$ $3267.00\pm 915.37$ $4.8$ $65.14$
Grams $7536.00\pm 240.38$ $1.41\pm 0.46$ $704.68\pm 123.61$ $1864.92\pm 204.72$ $4962.38\pm 1373.46$ $3.2$ $85.91$
Ano (Ours) $\mathbf{9556.00}\pm\mathbf{2370.99}$ $1.13\pm 0.43$ $\mathbf{733.95}\pm\mathbf{235.51}$ $2862.24\pm 365.46$ $\mathbf{5181.70}\pm\mathbf{1210.60}$ $\mathbf{1.8}$ $\mathbf{98.72}$
Table 5: Comparison of the average performance (± IC95%) of different optimizers across Atari environments.
As shown in Table 5, Ano and Adam are the strongest overall baselines. In the untuned setting, their mean ranks are 2.2 and 2.4; with tuning, 1.8 and 2.8, respectively. Ano attains the best average normalized score and the best mean rank in both regimes. A per-game breakdown indicates that Ano outperforms Adam on BattleZone, Name This Game, and Q*bert, whereas Adam clearly leads on Phoenix. For DoubleDunk, all optimizers (except Lion) saturate in the same plateau region (Fig. 4(b)), so no clear winner emerges. Interestingly, Adam performs slightly worse under our tuned configuration than under its default hyperparameters ($\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$). Our tuned settings were obtained via a grid search on MuJoCo with SAC and then transferred to Atari PPO (Appendix B), following the common practice that PPO and SAC often share learning-rate and momentum parameters. This transfer appears more benign for Ano benefiting from $\beta_{1}=0.95$, than for Adam, suggesting greater hyperparameter sensitivity of the latter.
@@ 7 Ablation Study
We conduct ablation studies on Ano and on its variant Anolog to quantify the contribution of each design component and, in particular, to justify adopting a logarithmic momentum schedule instead of a theoretically motivated square-root schedule.
Table 6 lists all ablated variants. We compared the performance across four benchmarks: deep reinforcement learning (HalfCheetah-v5), computer vision, and two NLP classification tasks—MRPC (relatively noisy) and SST‑2 (more stable).
Optimizer Base LR Adam Adapt. Yogi Adapt. Grad. Norm Mom. Norm Mom. Dir. Decoup. WD $\beta_{1,k}$ Score DRL Acc. (%) CIFAR-100 Acc. (%) MRPC Acc. (%) SST-2
@@
Adam $\beta_{1}$ $8048.02\pm 1323.36$ $69.73\pm 0.30$ $86.03\pm 1.66$ $93.08\pm 0.54$
Yogi $\beta_{1}$ $8871.78\pm 671.222$ $70.57\pm 0.07$ $85.05\pm 2.17$ $92.81\pm 0.42$
Grams $\beta_{1}$ $5493.76\pm 782.37$ $70.20\pm 0.17$ $81.94\pm 1.16$ $92.28\pm 0.40$
Yogi+Signum $\beta_{1}$ $260.91\pm 41.114$ $3.99\pm 2.01$ $68.38\pm 0.00$ $50.92\pm 0.00$
Signum+WeightDecay $\beta_{1}$ $8987.63\pm 1399.78$ $64.983\pm 1.59$ $86.03\pm 0.56$ $90.29\pm 0.27$
Signum+WeightDecay+GradNorm $\beta_{1}$ $53.63\pm 0.30$ $68.38\pm 0.00$ $53.33\pm 2.62$
Adam+GradNorm $\beta_{1}$ $9917.19\pm 1192.77$ $70.32\pm 0.64$ $86.52\pm 0.73$ $92.60\pm 1.01$
Ours (Ano) $\beta_{1}$ $\mathbf{10823.5}\pm\mathbf{416.07}$ $69.64\pm 0.79$ $86.44\pm 0.85$ $92.66\pm 0.45$
@@
Anoall 11/k1-1/k $212.18\pm 22.25$ $29.48\pm 2.4$ $76.42\pm 3.38$ $74.00\pm 11.97$
Anosqrt $1-1/\sqrt{k}$ $8674.09\pm 1390.00$ $67.26\pm 0.41$ $85.78\pm 0.96$ $91.44\pm 0.50$
Ours (Anolog) $1-1/\log k$ $\mathbf{9342.92}\pm\mathbf{1652.96}$ $67.00\pm 0.80$ $\mathbf{86.961}\pm\mathbf{1.181}$ $92.93\pm 0.15$
@@
As shown in Table 6, Ano achieves the highest mean return in deep reinforcement learning (10365.5 $\pm$ 513.5), representing an improvement of approximately 7% over the same algorithm using Adam-style second moments, and around 20% over Yogi and Signum, the strongest baselines. At the same time, Ano remains within 1% of the top accuracy on each supervised learning task. Performance degrades when either adaptive second-moment estimation or gradient normalization is removed, highlighting their complementary roles. Yogi-style second-moment updates outperform Adam-style ones in DRL and are therefore adopted. Regarding momentum schedules, the logarithmic schedule yields a 12% improvement in DRL return compared to the $\sqrt{k}$ schedule, without compromising classification accuracy, thus motivating its inclusion in the final design.

### v2

@@ 6.2 Natural Language Processing
All runs finetune the public bert-base-uncased checkpoint (devlin2019bert) with max sequence length 128, batch size 32, weight decay 0.01, linear schedule with 10% warmup, mixed precision, and 3 epochs (5 for small datasets: CoLA, MRPC, RTE, STS-B). Each configuration is repeated over 5 seeds.
@@
Default
Adam $\mathbf{59.40}\pm\mathbf{1.67}$ $\mathbf{84.62}\pm\mathbf{0.10}$ $88.06\pm 0.82$ $\mathbf{91.60}\pm\mathbf{0.15}$ $89.64\pm 0.10$ $66.67\pm 1.59$ $92.73\pm 0.46$ $88.44\pm 0.27$ $82.64$
@@
Table 3: Average performance (mean $\pm$ CI95%) of different optimizers on GLUE benchmark tasks (best per column in bold).
@@ 6.3 Deep Reinforcement learning
Reinforcement learning (RL) presents unique challenges, such as high gradient variance and the non-stationarity of the environment, both of which can strongly affect the behavior of optimization algorithms (henderson2018matters; franccois2018introduction). Because RL is inherently noisy and largely non-stationary, it is precisely the setting where Ano is expected to provide the largest benefits. For computational efficiency, we perform grid searches on HalfCheetah using 100k training steps. We acknowledge that this shorter horizon may bias tuning toward larger learning rates, which can be suboptimal for longer runs (e.g., 1M steps). To mitigate this, for each baseline optimizer we report the best performance between its default and tuned hyperparameters, ensuring that no method is disadvantaged by the tuning protocol. The tables indicate which configuration (default or tuned) was used, with full results provided in D.
@@
In this section, we employ the Soft Actor-Critic (SAC) algorithm (haarnoja2018soft) in the MuJoCo suite from the Gymnasium framework (todorov2012mujoco; towers2024gymnasiumstandardinterfacereinforcement). We reuse the standard SAC hyperparameter (full list in Appendix E.1 - Tab 11), as reported in the original work and subsequent studies and only vary the optimizer for actor, critics, and temperature. For each optimizer, we run 10 seeds, with 1M steps. We report below the average mean score on a 50-episodes test evaluation.
@@
Default
Adam $10549.48\pm 721.55$ $4336.64\pm 698.72$ $\mathbf{5357.14}\pm\mathbf{211.97}$ $4462.51\pm 588.77$ $3164.71\pm 600.48$ $3.4$ $90.66$
RMSprop $10506.23\pm 852.19$ $4234.37\pm 763.65$ $5395.51\pm 126.80$ $4160.06\pm 480.62$ $2973.86\pm 571.05$ $5.6$ $87.83$
Adan $7805.20\pm 1154.02$ $2985.19\pm 1018.79$ $5080.74\pm 305.26$ $4092.13\pm 379.92$ $3222.62\pm 235.25$ $5.6$ $78.38$
Lion $9527.96\pm 805.42$ $4948.26\pm 243.05$ $98.22\pm 32.33$ $4612.63\pm 367.77$ $3087.27\pm 628.06$ $4.4$ $71.74$
Grams $6782.60\pm 715.12$ $3207.30\pm 531.06$ $5104.10\pm 692.14$ $3656.66\pm 658.82$ $1475.34\pm 927.22$ $6.4$ $65.88$
Ano (Ours) $\mathbf{10864.09}\pm\mathbf{1052.24}$ $\mathbf{5285.44}\pm\mathbf{729.86}$ $5255.62\pm 815.92$ $\mathbf{5227.86}\pm\mathbf{436.49}$ $\mathbf{3535.32}\pm\mathbf{780.96}$ $\mathbf{1.4}$ $\mathbf{99.48}$
Anolog (Ours) $10557.05\pm 560.70$ $5089.12\pm 522.94$ $5242.78\pm 173.98$ $4606.02\pm 478.36$ $3314.12\pm 539.95$ $2.6$ $94.50$
Best Version
Adam [Default] $10549.48\pm 721.55$ $4336.64\pm 698.72$ $\mathbf{5357.14}\pm\mathbf{211.97}$ $4462.51\pm 588.77$ $3164.71\pm 600.48$ $4.6$ $90.38$
RMSprop [Default] $10506.23\pm 852.19$ $4234.37\pm 763.65$ $5395.51\pm 126.80$ $4160.06\pm 480.62$ $2973.86\pm 571.05$ $5.6$ $87.83$
Adan [Tuned] $10822.40\pm 475.75$ $5239.69\pm 270.96$ $4792.62\pm 904.44$ $4686.83\pm 502.28$ $3514.42\pm 143.57$ $3.2$ $95.01$
Lion [Tuned] $10482.06\pm 1018.86$ $4848.41\pm 821.79$ $1349.15\pm 1322.56$ $4876.76\pm 253.22$ $\mathbf{3592.87}\pm\mathbf{70.26}$ $4.2$ $81.30$
Grams [Tuned] $10533.70\pm 866.69$ $4607.59\pm 505.08$ $5147.04\pm 487.55$ $4644.45\pm 498.08$ $3147.82\pm 605.03$ $5.0$ $91.20$
Ano (Ours) [Default] $\mathbf{10864.09}\pm\mathbf{1052.24}$ $\mathbf{5285.44}\pm\mathbf{729.86}$ $5255.62\pm 815.92$ $\mathbf{5227.86}\pm\mathbf{436.49}$ $3535.32\pm 780.96$ $\mathbf{1.6}$ $\mathbf{99.16}$
Anolog (Ours) [Default] $10557.05\pm 560.70$ $5089.12\pm 522.94$ $5242.78\pm 173.98$ $4606.02\pm 478.36$ $3314.12\pm 539.95$ $2.6$ $94.20$
Table 4: Comparison of the IQM $\pm$ CI95% of different optimizers across environments.
@@
As summarized in Table 4, Ano performs favorably compared to Adam and other baselines across the MuJoCo tasks. On average, it achieves a +10% improvement in normalized score [^2] The normalized average is obtained by linearly rescaling each score between the minimum and maximum values observed across optimizers, followed by averaging. Complete normalized results are reported in Table 8., both under default and tuned hyperparameters. Without tuning, Ano ranks first in 4 out of 5 tasks; with best version, it remains the top optimizer in 3 out of 5 tasks. Although not always the best performer, Ano consistently ranks among the strongest optimizers, with its scores typically within or close to the 95% confidence intervals of the best baselines. Figure 2 shows that Ano reaches the final performance of Adam using approximately 50–70% fewer training steps, except for Humanoid. To address potential concerns about hyperparameter tuning, we evaluated the sensitivity to learning rate and momentum coefficients on a 100k-step HalfCheetah proxy (Figure 3; see B for full details). Ano shows lower sensitivity than Adam to both learning rate and betas, suggesting that its performance gains are not solely due to more favorable hyperparameter choices.
@@
To assess the generality of our findings, we additionally evaluate a discrete-action variant of Proximal Policy Optimization (PPO) (schulman2017proximalpolicyoptimizationalgorithms) on the Atari Learning Environment (ALE) (bellemare2013arcade). For compute efficiency, we report results on the Atari-5 subset proposed by Aitchison et al. (aitchison2023atari), which explains $98.4\%$ of the variance in full-suite performance ($R^{2}{=}0.984$). We use the reference PPO implementation from CleanRL (huang2022cleanrl) and keep its default network and optimization hyperparameters (full list in Appendix E.1 - tab 12). Environments are instantiated via EnvPool (weng2022envpoolhighlyparallelreinforcement) for fast batched simulation. Observations are resized to $84{\times}84$, converted to grayscale, and stacks of 4 consecutive frames are fed to the agent. We apply action repeat of 4, up to 30 no-op actions at reset, the ALE sticky-action protocol with repeat probability $0.25$ (machado2018revisiting), and FireReset where required. We used the full action set. During training, rewards are clipped to [1,1][-1,1]; evaluation uses unclipped rewards. We train for 1010M agent steps (with action repeat 4, i.e., $\approx 40$M ALE frames), evaluating every 200200k steps on 50 runs. Each checkpoint is evaluated with the same wrappers as training (except reward clipping). We report the average final score on the last evaluation. The normalization average is computed in the same ways that previous part [^3] We normalized each score by environment by the theoretically minimum (-18 for DoubleDunk, 0 for the others) and the maximum score obtained by the baselines, we then compute the mean average normalized score..
@@
Default
Adam $7615.00\pm 1299.89$ $1.08\pm 0.21$ $665.35\pm 64.89$ $\textbf{3443.30}\pm\textbf{240.68}$ $257.80\pm 135.90 4.4$4.4 $87.54$
RMSprop $7686.67\pm 859.43$ -$\textbf{-0.67}\pm\textbf{0.22}$ $798.00\pm 118.66$ $3031.13\pm 410.22$ $4585.67\pm 538.44$ $2.4$ $90.09$
Adan $6480.00\pm 702.14$ $0.91\pm 0.23$ $638.35\pm 18.08$ $2106.90\pm 110.74$ $2665.00\pm 708.69$ $5.4$ $74.11$
Lion $1392.00\pm 139.66$ $1.46\pm 0.65$ $508.15\pm 91.28$ $2432.35\pm 234.61$ $3768.00\pm 520.95$ $6.8$ $61.36$
Grams $7908.00\pm 687.61$ $0.93\pm 0.24$ $633.80\pm 23.52$ $2234.40\pm 130.73$ $3670.12\pm 634.75$ $4.6$ $82.41$
Ano (Ours) $\mathbf{8095.00}\pm\mathbf{494.70}$ $0.97\pm 0.14$ $\mathbf{845.35}\pm\mathbf{56.39}$ $2813.00\pm 243.04$ $\mathbf{4828.80}\pm\mathbf{386.73}$ $\mathbf{2.2}$ $\mathbf{95.99}$
Anolog (Ours) $7485.00\pm 1010.66$ $0.98\pm 0.14$ $751.05\pm 74.60$ $2983.00\pm 236.42$ $4773.75\pm 602.39$ $3.6$ $93.00$
Best Version
Adam [Baseline] $7615.00\pm 1299.89$ $1.08\pm 0.21$ $665.35\pm 64.89$ $\textbf{3443.30}\pm\textbf{240.68}$ $257.80\pm 135.90 4.4$4.4 $87.54$
RMSprop [Baseline] $7686.67\pm 859.43$ -$\textbf{-0.67}\pm\textbf{0.22}$ $798.00\pm 118.66$ $3031.13\pm 410.22$ $4585.67\pm 538.44$ $2.4$ $90.09$
Adan [Tuned] $4840.00\pm 2601.35$ $0.95\pm 0.23$ $754.20\pm 51.44$ $2647.20\pm 534.22$ $524.75\pm 448.74 4.4$4.4 $79.67$
Lion [Baseline] $1392.00\pm 139.66$ $1.46\pm 0.65$ $508.15\pm 91.28$ $2432.35\pm 234.61$ $3768.00\pm 520.95$ $6.8$ $61.36$
Grams [Tuned] $7715.00\pm 627.92$ $1.35\pm 0.75$ $690.40\pm 76.27$ $1989.15\pm 201.64$ $5049.25\pm 624.43$ $4.4$ $82.26$
Ano (Ours) [Tuned] $\mathbf{8625.00}\pm\mathbf{1870.44}$ $0.91\pm 0.21$ $\mathbf{828.10}\pm\mathbf{67.66}$ $2824.85\pm 226.30$ $\mathbf{5960.88}\pm\mathbf{912.36}$ $\mathbf{1.8}$ $\mathbf{96.13}$
Anolog (Ours) [Baseline] $7485.00\pm 1010.66$ $0.98\pm 0.14$ $751.05\pm 74.60$ $2983.00\pm 236.42$ $4773.75\pm 602.39$ $3.8$ $88.48$
Table 5: Comparison of the IQM $\pm$ CI95% of different optimizers across Atari environments.
As shown in Table 5, Ano and RMSprop perform strongest overall among the baselines. In the default setting, their mean ranks are 2.2 and 2.4; with the best version, 1.8 and 2.4, respectively. Ano achieves the highest average normalized score and mean rank in both regimes, with approximately 6–7% higher normalized average than RMSprop and 10% higher than Adam. Notably, Ano outperforms Adam on BattleZone, Name This Game, and Q*bert, whereas Adam and RMSprop perform best on Phoenix. For DoubleDunk, all optimizers (except Lion) plateau at similar levels (Fig. 4(b)), so no clear winner emerges.
@@ 7 Ablation Study
We conduct ablation studies on Ano and on its variant Anolog to quantify the contribution of each design component and, to justify adopting a logarithmic momentum schedule instead of a theoretically motivated square-root schedule.
We conduct ablation studies on Ano and its variant Anolog to quantify the contribution of each design component and to justify using a logarithmic momentum schedule rather than the theoretically motivated square-root schedule. Table 6 summarizes all ablated variants. To provide a comprehensive evaluation, we compare performance on four benchmarks: HalfCheetah from MuJoCo (todorov2012mujoco), CIFAR100 (krizhevsky2009learning), and two tasks from the GLUE benchmark: the small and noisy MRPC task, and the larger, more stable SST2 task. We follow the same experimental protocols as in Section 6, except that for HalfCheetah we train for 500k steps.
Optimizer Second Mom. Rule Grad. Norm Mom. Norm Mom. Dir. Decoup. WD $\beta_{1,k}$ Score DRL Acc. (%) CIFAR-100 Acc. (%) MRPC Acc. (%) SST-2
@@
Adam Adam $\beta_{1}$ $7480.55\pm 1323.36$ $69.84\pm 0.22$ $85.93\pm 0.92$ $\textbf{93.03}\pm\textbf{0.30}$
[…]
Grams Adam $\beta_{1}$ $5567.12\pm 782.37$ $70.20\pm 0.17$ $82.25\pm 0.74$ $92.29\pm 0.23$
[…]
Signum $\beta_{1}$ $9393.64\pm 1399.78$ $65.11\pm 0.90$ $86.42\pm 0.72$ $90.41\pm 0.30$
SignumGrad $\beta_{1}$ $53.93\pm 0.68$ $68.38\pm 0.00$ $53.33\pm 2.62$
AdamGrad Adam $\beta_{1}$ $9855.19\pm 1173.19$ $70.30\pm 0.38$ $86.96\pm 0.85$ $92.71\pm 0.45$
AnoWoTweak Yogi $\beta_{1}$ $9053.10\pm 792.13$ $\textbf{70.32}\pm\textbf{1.20}$ $\textbf{87.06}\pm\textbf{0.69}$ $92.80\pm 0.45$
[…]
@@
[…]
@@
As shown in Table 6, Ano achieves the highest mean return in deep reinforcement learning, improving by roughly 7% over the same algorithm with Adam-style second moments and about 15% over Ano with Yogi-style second moments, while staying within 1% of the best accuracy on all supervised learning tasks. Using only the sign of the momentum (e.g., Signum, AdamGrad, AnoWoDecay) also improves DRL performance, supporting our design choice to decouple sign and magnitude: this enables larger update steps, which are particularly beneficial in noisy or non-stationary environments. Performance drops when either gradient normalization (SignumGrad) or gradient magnitude (YogiSignum) is removed, underscoring their complementary roles. For momentum schedules, the logarithmic schedule improves DRL return over the $\sqrt{k}$ schedule while staying within the 95% confidence interval on other tasks, motivating its inclusion in the final design.